cos(3t), t<n/2, 2. Let f(t) = sin(2t), 7/2<t< , Write f(t) in terms of the unit step e3 St. function. Then find c{f(t)}.
Answer #1

We have formula if f(t) =( a b c iftat, iftistata if tast then f(t) = at (b-a) Ult-t) + (6-b) Ult- tz) Put a = cos(3+), be si541633€) + 16+sin(at) –sin (36) ? thom et sin (24)) ets TS (ata) HD S-3